# Inner Thinking Transformer: Leveraging Dynamic Depth Scaling to Foster Adaptive Internal Thinking

### v1

@@ 1 Introduction
Large language models (LLMs) Anthropic (2023); OpenAI (2023); Touvron et al. (2023) have demonstrated remarkable performance across numerous natural language tasks. Recent studies Fernandez et al. (2024); Hoffmann et al. (2022); Chen et al. (2024) indicate that scaling laws for LLM parameters exhibit diminishing returns under constrained data availability and computational resource budgets. Scaling model parameters increases computational and deployment costs, making high-performance models impractical for resource-constrained environments. Meanwhile, smaller models encounter performance bottlenecks primarily attributable to limited parameter space.
Recent approaches, such as Test-Time Scaling ("Slow-Thinking") Muennighoff et al. (2025); Snell et al. (2024); Ma et al. (2024), aim to enhance performance by allocating more computation during the inference search process. While effective, these methods are limited by the reliance on accurately generating key tokens, which can lead to catastrophic reasoning failures Chen et al. (2023a); Singh et al. (2024); Jiang et al. (2024b), especially in smaller models. Some works enhance model performance through layer sharing Li et al. (2024b, c), recursion Ng and Wang (2024); Dehghani et al. (2019b); Geiping et al. (2025), or implicit reasoning Deng et al. (2023); Shalev et al. (2024), but they fail to flexibly improve the model’s reasoning ability on key tokens, which either suffer from insufficient performance or redundant overhead.
@@ 2 Observation
To investigate how models learn about critical tokens, our empirical analysis of GPT-2’s attention matrices through gradient nuclear norm (GNN) measurements Li et al. (2024a) reveals systematic patterns in layer-wise dynamics. Using the AQuA corpus Ling et al. (2017), we firstly train GPT-2 in 100 samples then categorize samples in evaluation into easy (model answers correctly) and hard (model answers incorrectly). In Figure 2, for easy samples, GNN values decay exponentially across early layers (L0-L2) and final layers (L11), stabilizing below 3 in layers (L3-L10). In contrast, hard samples exhibit persistent GNN oscillations throughout all 12 layers, punctuated by abrupt spikes at strategic layer positions (L3, L5, L7, L9).
@@ 3.2 Residual Thinking Connection
where $\phi^{(t)}\in\mathbb{R}^{d}$ the learnable thinking position encoding associated with the $t$-th inner thinking step, which measuring the differences and importance of each step. Rather than processing the input representation only once, RTC Layer iteratively refine it by adding the residual contributions of each step’s layer-output together with a learnable encoding. Compared to direct looping Ng and Wang (2024); Dehghani et al. (2019b), RTC not only enables deeper thinking but also effectively measures and combines each thinking step, allowing them to complement each other. RTC provides the foundation for scaling Inner Thinking during testing.
@@ 4.1 Setup
  • The Loop Neural Network design Ng and Wang (2024); Dehghani et al. (2019b); Geiping et al. (2025) implements model-level recurrence for iterative refinement.
@@ 5 Related Work
. The concept of recurrence in machine learning traces back to foundational works on neural computation (Braitenberg, 1986) and LSTM networks (Gers and Schmidhuber, 2000). Modern extensions integrate recurrence into transformers through depth recurrence (Dehghani et al., 2019a; Lan et al., 2019; Ng and Wang, 2024). Recent works have re-discovered this idea for implicit reasoning Deng et al. (2023); Hao et al. (2024) and test-time scaling Geiping et al. (2025). In contrast, ITT establishes a general-purpose recursive reasoning framework within individual layers and designs the Residual Thinking Cnnection (RTC) for enhanced capability.
@@
, like Mixture-of-Expert (MoE), reduce computational overhead by activating only a subset of networks Fedus et al. (2022); Riquelme et al. (2021); Zhou et al. (2022); Jiang et al. (2024a); Xue et al. (2024). Some works focus on elastic computation in depth, such as early exit Elhoushi et al. (2024); Chen et al. (2023b), parameter sharing Li et al. (2024b, c) or using token-routing for dynamic layer skipping Zhang et al. (2024). Inspired by these works, ITT designs an elastic deep thinking architecture with Adaptive Token Routing (ATR) for efficient and adaptive computational resources allocation.
@@ Acknowledgments
We would like to thank members of the IIE KDsec group for their valuable feedback and discussions. We are very grateful to Mengzhou Xia for providing the concise and effective ShearingLLaMA experimental code and for her assistance during the reproduction process. Work done during Yilong Chen’s internship in Baidu Inc. This research is supported by the Youth Innovation Promotion Association of CAS (Grant No.2021153).
@@ A.2 Extend Related Work
The concept of recurrence in machine learning traces back to foundational works on neural computation (Braitenberg, 1986) and LSTM networks (Gers and Schmidhuber, 2000). Modern extensions integrate recurrence into transformers through depth recurrence (Dehghani et al., 2019a; Lan et al., 2019; Ng and Wang, 2024), with recent improvements demonstrating algorithmic generalization via randomized unrolling (Schwarzschild et al., 2021; McLeish et al., 2024). From an optimization perspective, these models relate to energy-based gradient dynamics (LeCun et al., 2006) and test-time adaptation (Boudiaf et al., 2022). Recent works have introduced it for implicit reasoning Deng et al. (2023); Hao et al. (2024) and test-time scaling Geiping et al. (2025). Inspired by these, ITT focuses on recursive reasoning within individual layers and designs the RTC architecture with theoretical support to enhance this capability.
@@
Dynamic Computation Allocation in architectures, like Sparse Mixture-of-Expert (MoE), utilize input adaptivity to reduce computational overhead by activating only a subset of subnetworks, or "experts," for each input token Fedus et al. (2022); Riquelme et al. (2021); Zhou et al. (2022); Jiang et al. (2024a); Xue et al. (2024). Recent developments have introduced heterogeneous experts, integrating experts with varying capacities and specializations Wu et al. (2024); He (2024); Dean (2021); Zhou et al. (2022). Some works focus on elastic computation in depth, such as early exit Elhoushi et al. (2024); Chen et al. (2023b), parameter sharing Li et al. (2024b, c) or using token-routing for dynamic layer skipping (Mixture of Depth) Zhang et al. (2024). Inspired by these works, ITT designs an elastic deep thinking architecture and uses Residual Thinking Connections to address the issue of non-continuous layer skipping.

### v2

@@ 1 Introduction
Large language models (LLMs) Anthropic (2023); OpenAI (2023); Touvron et al. (2023) have demonstrated remarkable performance across numerous natural language tasks. Recent studies Fernandez et al. (2024); Hoffmann et al. (2022); Wang et al. (2024b) indicate that scaling laws for LLM parameters exhibit diminishing returns under constrained data availability and computational resource budgets. Scaling model parameters increases computational and deployment costs, making high-performance models impractical for resource-constrained environments. Meanwhile, smaller models encounter performance bottlenecks primarily attributable to limited parameter space.
Recent approaches, such as Test-Time Scaling ("Slow-Thinking") Muennighoff et al. (2025); Snell et al. (2024); Ma et al. (2025), aim to enhance performance by allocating more computation during the inference search process. While effective, these methods are limited by the reliance on accurately generating key tokens, which can lead to catastrophic reasoning failures Lin et al. (2025); Singh et al. (2024); Jiang et al. (2024b), especially in smaller models. Some works enhance model performance through layer sharing Mu et al. (2024); Wang et al. (2024a), recursion Ng and Wang (2024); Dehghani et al. (2019a); Geiping et al. (2025), or implicit reasoning Deng et al. (2023); Shalev et al. (2024), but they fail to flexibly improve the model’s reasoning ability on key tokens, which either suffer from insufficient performance or redundant overhead.
@@ 2 Observation
To investigate how models learn about critical tokens, our empirical analysis of GPT-2’s attention matrices through gradient nuclear norm (GNN) measurements Li et al. (2024) reveals systematic patterns in layer-wise dynamics. Using the AQuA corpus Ling et al. (2017), we firstly train GPT-2 in 100 samples then categorize samples in evaluation into easy (model answers correctly) and hard (model answers incorrectly). In Figure 2, for easy samples, GNN values decay exponentially across early layers (L0-L2) and final layers (L11), stabilizing below 3 in layers (L3-L10). In contrast, hard samples exhibit persistent GNN oscillations throughout all 12 layers, punctuated by abrupt spikes at strategic layer positions (L3, L5, L7, L9).
@@ 3.2 Residual Thinking Connection
where $\phi^{(t)}\in\mathbb{R}^{d}$ the learnable thinking position encoding associated with the $t$-th inner thinking step, which measuring the differences and importance of each step. Rather than processing the input representation only once, RTC Layer iteratively refine it by adding the residual contributions of each step’s layer-output together with a learnable encoding. Compared to direct looping Ng and Wang (2024); Dehghani et al. (2019a), RTC not only enables deeper thinking but also effectively measures and combines each thinking step, allowing them to complement each other. RTC provides the foundation for scaling Inner Thinking during testing.
@@ 4.1 Setup
  • The Loop Neural Network design Ng and Wang (2024); Dehghani et al. (2019a) implements recurrence for iterative refinement.
@@ 5 Related Work
The concept of recurrence in machine learning traces back to foundational works on neural computation (Braitenberg, 1986) and LSTM networks (Gers et al., 2000). Modern extensions integrate recurrence into transformers through depth recurrence (Dehghani et al., 2019b; Lan et al., 2020; Ng and Wang, 2024). Recent works have re-discovered this idea for implicit reasoning Deng et al. (2023); Hao et al. (2024) and test-time scaling Geiping et al. (2025). In contrast, ITT establishes a general-purpose recursive reasoning framework within individual layers and designs the Residual Thinking Cnnection (RTC) for enhanced capability.
@@
Dynamic Computation Allocation, like Mixture-of-Expert (MoE), reduce computational overhead by activating only a subset of networks Fedus et al. (2022); Riquelme et al. (2021); Zhou et al. (2022); Jiang et al. (2024a); Xue et al. (2024). Some works focus on elastic computation in depth, such as early exit Elhoushi et al. (2024); Chen et al. (2024), parameter sharing Mu et al. (2024); Wang et al. (2024a) or using token-routing for dynamic layer skipping Zhang et al. (2024). Inspired by these works, ITT designs an elastic deep thinking architecture with Adaptive Token Routing (ATR) for efficient and adaptive computational resources allocation.
@@ Acknowledgments
We would like to thank members of the IIE KDsec group for their valuable feedback and discussions. We sincerely thank Sean McLeish for his diligent review and critical feedback on this work. We are very grateful to Mengzhou Xia for providing the concise and effective ShearingLLaMA experimental code and for her assistance during the reproduction process. Work done during Yilong Chen’s internship in Baidu Inc. This research is supported by the Youth Innovation Promotion Association of CAS (Grant No.2021153).
@@ A.2 Extend Related Work
The concept of recurrence in machine learning traces back to foundational works on neural computation (Braitenberg, 1986) and LSTM networks (Gers et al., 2000). Modern extensions integrate recurrence into transformers through depth recurrence (Dehghani et al., 2019b; Lan et al., 2020; Ng and Wang, 2024), with recent improvements demonstrating algorithmic generalization via randomized unrolling (Schwarzschild, 2023; Tu et al., 2020). From an optimization perspective, these models relate to energy-based gradient dynamics (Hinton et al., 2006) and test-time adaptation (Xiao and Snoek, 2024). Recent works have introduced it for implicit reasoning Deng et al. (2023); Hao et al. (2024) and test-time scaling Geiping et al. (2025). Inspired by these, ITT focuses on recursive reasoning within individual layers and designs the RTC architecture with theoretical support to enhance this capability.
@@
Dynamic Computation Allocation in architectures, like Sparse Mixture-of-Expert (MoE), utilize input adaptivity to reduce computational overhead by activating only a subset of subnetworks, or "experts," for each input token Fedus et al. (2022); Riquelme et al. (2021); Zhou et al. (2022); Jiang et al. (2024a); Xue et al. (2024). Recent developments have introduced heterogeneous experts, integrating experts with varying capacities and specializations Wu et al. (2024); He (2024); Dean (2021); Zhou et al. (2022). Some works focus on elastic computation in depth, such as early exit Elhoushi et al. (2024); Chen et al. (2024), parameter sharing Mu et al. (2024); Wang et al. (2024a) or using token-routing for dynamic layer skipping (Mixture of Depth) Zhang et al. (2024). Inspired by these works, ITT designs an elastic deep thinking architecture and uses Residual Thinking Connections to address the issue of non-continuous layer skipping.